# A Generative Approach to Control Complex Physical Systems

### v1

@@ Abstract
Controlling the evolution of complex physical systems is a fundamental task across science and engineering. Classical techniques suffer from limited applicability or huge computational costs. On the other hand, recent deep learning and reinforcement learning-based approaches often struggle to optimize long-term control sequences under the constraints of system dynamics. In this work, we introduce Diffusion Physical systems Control (DiffPhyCon), a new class of method to address the physical systems control problem. DiffPhyCon excels by simultaneously minimizing both the learned generative energy function and the predefined control objectives across the entire trajectory and control sequence. Thus, it can explore globally and identify near-optimal control sequences. Moreover, we enhance DiffPhyCon with prior reweighting, enabling the discovery of control sequences that significantly deviate from the training distribution. We test our method in 1D Burgers’ equation and 2D jellyfish movement control in a fluid environment. Our method outperforms widely applied classical approaches and state-of-the-art deep learning and reinforcement learning methods. Notably, DiffPhyCon unveils an intriguing fast-close-slow-open pattern observed in the jellyfish, aligning with established findings in the field of fluid dynamics.
@@ 1 Introduction
Modeling the dynamics of complex physical systems is an important class of problems in science and engineering. Usually, we are not only interested in predicting a physical system’s behavior but also injecting time-variant signals to steer its evolution and optimize specific objectives. This gives rise to the complex physical control problem, a fundamental task with numerous applications, including controlled nuclear fusion [9], fluid control [21], underwater devices [72] and aviation [48], among others.
@@
We demonstrate the effectiveness of DiffPhyCon via extensive experiments on 1D Burgers’ equation and 2D jellyfish movement control tasks, with the additional challenge of partial observation and partial control. In particular, the 2D jellyfish movement control problem is an important source of inspiration for the design of underwater and aerial devices, and its challenges come from complex vortice behavior and fluid-solid coupling dynamics [55, 27]. We generate such a dataset based on a CFD (computational fluid dynamic) software to mimic the movement of a jellyfish under control signals of its flapping behavior. To advance research in controlling complex physical systems, we contribute our dataset as a benchmark. On both tasks, our method outperforms widely applied classical control methods and is also competitive with recent supervised learning and strong reinforcement learning baselines. Notably, DiffPhyCon reveals the intriguing fast-close-slow-open pattern exhibited by the jellyfish, aligning with findings in the field of fluid dynamics [27].
In summary, we contribute the following: (1) We develop DiffPhyCon, a novel generative method to control complex physical systems. By optimizing trajectory and control sequences jointly in the entire horizon by diffusion models, DiffPhyCon facilitates global optimization of long-term dynamics and helps to reduce myopic failure modes. (2) We introduce the prior reweighting technique to generate control sequences that are superior to those in the training set. (3) We demonstrate the effectiveness of our method on 1D Burgers’ equation and 2D jellyfish movement control tasks, particularly demonstrating advantages in scenarios with partial observations/control. (4) We release a new benchmark of jellyfish movement control to the community of complex physical system control.
@@ 2.2 Preliminary: Diffusion Models
Notation. We use $\mathbf{v}_{[n,m]}=[\mathbf{v}_{n},\cdots,\mathbf{v}_{m}]$ to denote a sequence of variables. We use $\mathbf{z}_{[0,T-1],k}$ to denote the hidden variable of $\mathbf{z}_{[0,T-1]}$ in a diffusion step $k$. For simplicity, we abbreviate $\mathbf{w}_{[0,T-1]}$, $\mathbf{u}_{[1,T]}$ as $\mathbf{w}$, $\mathbf{u}$. Concatenation of two variables are denoted via e.g., $[\mathbf{u},\mathbf{w}]$.
@@ 3.1 Generative Control by Diffusion Models
where $\lambda$ is a hyperparameter. This formulation optimizes $\mathbf{u}$ and $\mathbf{w}$ of all physical time steps simultaneously. The first term encourages the generated $\mathbf{w}$ and $\mathbf{u}$ to satisfy the PDE constraints $\mathcal{C}(\mathbf{u},\mathbf{w})=0$. The second term guides optimization towards optimal objectives. The advantage of this optimization framework is that it tasks a global optimization on $\mathbf{u}$ and $\mathbf{w}$ of all times steps, which may obtain better solutions that are faithful to the dynamics of the physical system. Details about the effect of the hyperparameter $\lambda$ are provided in Appendix J.
@@
$\hat{\mathbf{z}}_{k}=(\mathbf{z}_{k}-\sqrt{1-\bar{\alpha}_{k}}\bm{\epsilon}_{\theta}(\mathbf{z}_{k},\mathbf{c},k))/\sqrt{\bar{\alpha}_{k}},$ (6)
@@ 3.2 Prior Reweighting
$\nabla E_{\phi,\theta}^{(\gamma)}(\mathbf{u},\mathbf{w},\mathbf{c})=(\gamma-1)\nabla E^{(p)}_{\phi}(\mathbf{w},\mathbf{c})+\nabla E_{\theta}(\mathbf{u},\mathbf{w},\mathbf{c}),$ (12)
@@
iteratively, where $\mathbf{z}_{k}=[\mathbf{u}_{k},\mathbf{w}_{k}]$. The difference between this iteration scheme and Eq. (7) is that it uses an additional step to update $\mathbf{w}_{k}$ based on the predicted noise of $\bm{\epsilon}_{\phi}$. This guides $\mathbf{z}_{k}=[\mathbf{u}_{k},\mathbf{w}_{k}]$ to move towards the reweighted distribution $p_{\gamma}(\mathbf{u},\mathbf{w}|\mathbf{c})$ while aligning with the direction to decrease the objective by its guidance in the iteration of $\mathbf{z}_{k}$. The complete algorithm is present in Algorithm 1. Detailed discussion and results about how to set the hyperparameter $\gamma$ are provided in Appendix J.
@@ 4 Experiments
In this section, we aim to answer the following questions: (1) Can DiffPhyCon present superiority over traditional, supervised learning, and reinforcement learning methods for physical systems control? (2) Does the proposed prior reweighting technique help achieve better control objectives? (3) Could answers to (1) and (2) be generalized to more challenging partial observation or partial control scenarios? To answer these questions, we conduct experiments on the 1D Burgers’ equation and 2D jellyfish movement control problem, both of which are vital and challenging.
The following state-of-the-art control methods are selected as baselines. For 1D Burgers’ equation, we use (1) the classical and widely used control algorithm Proportional-Integral-Derivative (PID) [33] interacting with our trained surrogate model of the solver; (2) Supervised Learning method (SL) [23]; RL methods including (3) Soft Actor-Critic (SAC) [15] with offline and surrogate-solver versions; (4) Behaviour Cloning (BC) [50]; and (5) Behavior Proximal Policy Optimization (BPPO) [75]. Specifically, the surrogate-solver version of SAC interacts with our trained surrogate model of the solver, while the offline version only uses given data. BC and BPPO are also in offline version. For 2D jellyfish movement control, baselines include SL, SAC (offline), SAC (surrogate-solver), BC, BPPO, and an additional classical multi-input multi-output algorithm Model Predictive Control (MPC) [59]. PID is inapplicable to this data-driven task. Detailed descriptions of baselines are provided in Appendix F and Appendix G. We provide the code to reproduce the result for our method and baselines here.
@@
$\mathcal{J}_{\text{actual}}\coloneqq\int_{\Omega}|u(T,x)-u_{d}(x)|^{2}\mathrm{d}x,\ \ \mathcal{J}_{\text{energy}}\coloneqq\int_{[0,T]\times\Omega}|w(t,x)|^{2}\mathrm{d}t\mathrm{d}x$ (17)
To make the evaluation challenging, we select three different experiment settings that correspond to different real-life scenarios: partial observation, full control (PO-FC), full observation, partial control (FO-PC), and partial observation, partial control (PO-PC), which are elaborated on in Appendix C.2 and illustrated in Appendix B. These settings are challenging for classical control methods such as PID since they require capturing the long-range dependencies in the system dynamics. Note that the reported metrics in different settings are not directly comparable.
In this experiment, DiffPhyCon uses the guidance conditioning (Section 2) to optimize $\mathcal{J}_{\text{actual}}$ and explicit guidance to optimize the energy cost, which is elaborated in Appendix C.
@@
To compare the ability of different methods to optimize $\mathcal{J}_{\text{actual}}$ with constrained energy cost $\mathcal{J}_{\text{energy}}$, we compare the Pareto frontiers of different methods in Figure 3. We vary the hyperparameter $\lambda$ to control the tradeoff between $\mathcal{J}_{\text{actual}}$ and the energy cost since most baselines have this hyperparameter. As can be observed in Figure 3, the Pareto frontiers of DiffPhyCon are consistently among the best, achieving the lowest $\mathcal{J}_{\text{actual}}$ for most settings of the energy budget. Although SL performs well in full observation setting (b) where the system dynamics can be more easily predicted, it encounters difficulty in partial observation scenarios (a)(c). The results demonstrate DiffPhyCon’s ability to generate near-optimal control sequences compared to baselines. More visualization results are provided in Appendix B. For efficiency evaluation, please refer to Table 19 in Appendix I.
@@
Experiment settings. This task is to control the movement of a flapping jellyfish with two wings in a 2D fluid field where fluid flows at a constant speed. The jellyfish is propelled by the fluid when its wings flap. Its moving speed and efficiency are determined by the mode of flapping. This task represents a common phenomenon in aquatic locomotion, which involves complex fluid-solid interactions and is an important source of inspiration for the design of underwater and aerial devices [27]. For this task, fluid dynamics follows the 2D incompressible Navier-Stokes Equation:
@@
subject to Eq. (18) and the boundary condition that the velocity of fluid vanishes near the boundary. The hyperparameter $\zeta$ is set to be 1000. We evaluate in two settings: full observation, where the full state $\mathbf{u}=(\mathbf{v},p)$ is observed; and partial observation, where only pressure is observed. This task is very challenging due to the complicated dynamics of fluid-solid interactions and vortices behaviour [55, 27], especially in the scenario of partial observation where the missing of $\mathbf{v}$ in Eq. (18) restricts the information available to generate well-informed control signals. Details of the experiment are provided in Appendix E.
Open-source Dataset Description. We use the Lily-Pad simulator [68] to generate a dataset describing the movement of jellyfish under the control of its flapping behavior. Statistics about the dataset are listed in Table 2. The feature of this dataset is that it contains a rich variation of vortices behavior and fluid-solid interaction dynamics, determined by violent changes in opening angles of wings and open-close phase ratio. It would serve as an important benchmark for studying complex physical system control problems. This dataset is released via an anonymous link. Details about the dataset are provided in Appendix D.
@@
Results. Evaluation results are presented in Table 3. We see that our method outperforms the baselines by a large margin in optimizing the control objective $\mathcal{J}$. At a cost of slightly increasing the control cost $R(\mathbf{w})$, control sequences generated by our method achieve a much faster average speed than baselines. In the full observation setting, the control objective achieved by DiffPhyCon is -53.18 lower than the best baseline BPPO, while the average speed exhibits an increment of 172.2 over it. This demonstrates that diffusion models are effective for this challenging control task by performing global optimization of trajectory and control sequences in a generative approach. Comparison between DiffPhyCon-lite and DiffPhyCon reveals that by flattening the prior distribution of control sequences, another significant improvement is further achieved. Configuration of the hyperparameter $\gamma$ in DiffPhyCon and performance with respect to varying $\gamma$ is presented in Figure 14 in Appendix J. Even in the more challenging partial observation setting, DiffPhyCon still exhibits substantial advantages over existing methods. This reflects our method has a strong control capability under inadequate information.
@@
Figure 5: Movement and fluid field visualization on the jellyfish controlled by DiffPhyCon as in the middle subfigure of Figure 4.
@@
Figure 4 visualizes generated opening angle curves of different methods on three test jellyfish. Opening angle curves of DiffPhyCon-lite show an obvious fast-close-slow-open shape, which is proven to produce high speed in jellyfish movement [27]. The reason is that the fast closing of wings leads to a thrust from fluid in the early state, resulting in a long-term high speed, and followed by a slow opening to reduce resistance. While this mode of movement appears rarely in the training dataset, DiffPhyCon-lite could generate such control sequences for most test samples. This reflects that diffusion models under guidance are effective in optimizing the control objective. Furthermore, DiffPhyCon makes this mode of movement more aggressive, with sharper change of the opening angle in the beginning and end stage within a period. This provides strong evidence that reweighting the prior distribution of control sequences allows for flexible sampling over a flattened distribution, thus control sequences with low prior but good objectives are more likely to be sampled. The movement and the resulting fluid field of the jellyfish corresponding to the middle subfigure of Figure 4 controlled by DiffPhyCon is illustrated in Figure 5 and more examples are provided in Figure 13 in Appendix H.6. Conversely, opening angles obtained by baselines are inferior. The reason may be that they predict opening angles sequentially and hard to globally optimize the objective with three conflicting terms: average speed, $R(\mathbf{w})$, and $d(\mathbf{w}_{T},\mathbf{w}_{0})$. We further study such myopic failure mode of SAC in Appendix H.7. More comparisons about the variation of the weight $\zeta$ are listed in Table 16 and Table 17 in Appendix H.6. For efficiency evaluation, please refer to Table 20 in Appendix I.
We also extend the experiment setting to a high-dimensional control signal setting, where the wings of the jellyfish are assumed to be soft. We find that our method is still competitive with baselines. This demonstrates the good scalability of our method when facing higher-dimensional physical systems. Details about this evaluation are provided in Appendix K.
5 Limitation
There are still several limitations of our DiffPhyCon that provide exciting opportunity for future works. Firstly, our approach is data-driven, so it is not guaranteed to achieve optimal solutions and also lacks estimation of the error gap between the generated control sequence and the optimal ones. Secondly, the training of DiffPhyCon is currently conducted in an offline fashion, without interaction with a ground-truth solver. Incorporating solvers into the training framework could facilitate real-time feedback, enabling the model to adapt dynamically to the environment and discover novel strategies and solutions. Furthermore, our proposed DiffPhyCon presently operates in an open-loop manner, as it does not consider real-time feedback from solvers. Integrating such feedback would empower the algorithm to adjust its control decisions based on the evolving state of the environment.

### v3

@@ Abstract
Controlling the evolution of complex physical systems is a fundamental task across science and engineering. Classical techniques suffer from limited applicability or huge computational costs. On the other hand, recent deep learning and reinforcement learning-based approaches often struggle to optimize long-term control sequences under the constraints of system dynamics. In this work, we introduce Diffusion Physical systems Control (DiffPhyCon), a new class of method to address the physical systems control problem. DiffPhyCon excels by simultaneously minimizing both the learned generative energy function and the predefined control objectives across the entire trajectory and control sequence. Thus, it can explore globally and plan near-optimal control sequences. Moreover, we enhance DiffPhyCon with prior reweighting, enabling the discovery of control sequences that significantly deviate from the training distribution. We test our method on three tasks: 1D Burgers’ equation, 2D jellyfish movement control, and 2D high-dimensional smoke control, where our generated jellyfish dataset is released as a benchmark for complex physical system control research. Our method outperforms widely applied classical approaches and state-of-the-art deep learning and reinforcement learning methods. Notably, DiffPhyCon unveils an intriguing fast-close-slow-open pattern observed in the jellyfish, aligning with established findings in the field of fluid dynamics. The project website, jellyfish dataset, and code can be found at https://github.com/AI4Science-WestlakeU/diffphycon.
[…]
@@ 1 Introduction
Modeling the dynamics of complex physical systems is an important class of problems in science and engineering. Usually, we are not only interested in predicting a physical system’s behavior but also injecting time-variant signals to steer its evolution and optimize specific objectives. This gives rise to the complex physical control problem, a fundamental task with wide applications, such as controlled nuclear fusion [10], fluid control [22], underwater devices [70] and aviation [47], among others.
@@
In summary, we contribute the following: (1) We develop DiffPhyCon, a novel generative method to control complex physical systems. By optimizing trajectory and control sequences jointly in the entire horizon by diffusion models, DiffPhyCon facilitates global optimization of long-term dynamics and helps to reduce myopic failure modes. (2) We introduce the prior reweighting technique to plan control sequences that are superior to those in the training set. (3) We demonstrate the effectiveness of our method on 1D Burgers’ equation, 2D jellyfish movement control, and 2D high-dimensional smoke control tasks. On all three tasks, our method outperforms widely applied classical control methods and is also competitive with recent supervised learning and strong reinforcement learning baselines, particularly demonstrating advantages of DiffPhyCon in scenarios with partial observations and partial/indirect control. Notably, DiffPhyCon reveals the intriguing fast-close-slow-open pattern exhibited by the jellyfish, aligning with findings in the field of fluid dynamics [28]. (4) We generate a dataset of 2D jellyfish based on a CFD (computational fluid dynamic) software to mimic the movement of a jellyfish under control signals of its flapping behavior. To advance research in controlling complex physical systems, we release our dataset as a benchmark.
@@ 2.2 Preliminary: Diffusion Models
Notation. We use $\mathbf{v}_{[n,m]}=[\mathbf{v}_{n},\cdots,\mathbf{v}_{m}]$ to denote a sequence of variables. We use $\mathbf{z}_{[0,T-1],k}$ to denote the hidden variable of $\mathbf{z}_{[0,T-1]}$ in a diffusion step $k$. For simplicity, we abbreviate $\mathbf{w}_{[0,T-1]}$, $\mathbf{u}_{[1,T]}$ as $\mathbf{w}$, $\mathbf{u}$. Concatenation of two variables is denoted via e.g., $[\mathbf{u},\mathbf{w}]$.
@@ 3.1 Generative Control by Diffusion Models
where $\lambda$ is a hyperparameter. This formulation optimizes $\mathbf{u}$ and $\mathbf{w}$ of all physical time steps simultaneously. The first term encourages the generated $\mathbf{w}$ and $\mathbf{u}$ to satisfy the PDE constraints $\mathcal{C}(\mathbf{u},\mathbf{w})=0$. The second term guides optimization towards optimal objectives. The advantage of this optimization framework is that it tasks a global optimization on $\mathbf{u}$ and $\mathbf{w}$ of all times steps, which may obtain better solutions that are faithful to the dynamics of the physical system. Details about the effect of the hyperparameter $\lambda$ are provided in Appendix L.
@@
$\hat{\mathbf{z}}_{k}=(\mathbf{z}_{k}-\sqrt{1-\bar{\alpha}_{k}}\bm{\epsilon}_{\theta}(\mathbf{z}_{k},\mathbf{c},k))/\sqrt{\bar{\alpha}_{k}}.$ (6)
@@ 3.2 Prior Reweighting
$\nabla E_{\phi,\theta}^{(\gamma)}(\mathbf{u},\mathbf{w},\mathbf{c})$ $=(\gamma-1)\nabla E^{(p)}_{\phi}(\mathbf{w},\mathbf{c})$
$+\nabla E_{\theta}(\mathbf{u},\mathbf{w},\mathbf{c}),$
@@
iteratively, where $\mathbf{z}_{k}=[\mathbf{u}_{k},\mathbf{w}_{k}]$. The difference between this iteration scheme and Eq. (7) is that it uses an additional step to update $\mathbf{w}_{k}$ based on the predicted noise of $\bm{\epsilon}_{\phi}$. This guides $\mathbf{z}_{k}=[\mathbf{u}_{k},\mathbf{w}_{k}]$ to move towards the reweighted distribution $p_{\gamma}(\mathbf{u},\mathbf{w}|\mathbf{c})$ while aligning with the direction to decrease the objective by its guidance in the iteration of $\mathbf{z}_{k}$. The complete algorithm is present in Algorithm 1. Detailed discussion and results about how to set the hyperparameter $\gamma$ are provided in Appendix L.
Theoretical Analysis. Consider a pair $[\mathbf{u},\mathbf{w}]$ sampled using the prior reweighting technique with hyperparameter $\gamma$: $[\mathbf{u},\mathbf{w}]\sim p_{\gamma}(\mathbf{u},\mathbf{w})=p(\mathbf{u}|\mathbf{w})p^{\gamma}(\mathbf{w})/C_{\gamma}$. Denote $\mathcal{J}^{*}$ as the global minimum of the control objective $\mathcal{J}$. Define $Q(\varepsilon)$ to be the "$\epsilon$-optimal" solution set of $[\mathbf{u},\mathbf{w}]$ such that $\mathcal{J}(\mathbf{u},\mathbf{w})-\mathcal{J}^{*}\leq(\epsilon)$, whose complement set is $Q(\varepsilon)^{c}$, and denote $\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})$ as its indicator function, i.e. $\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})=1$ if $[\mathbf{u},\mathbf{w}]\in Q(\varepsilon)$; otherwise 0. Define $Y$ to be the random variable of “whether use $\mathcal{J}$ as a guidance for sampling”, namely,
$p(Y|\mathbf{u},\mathbf{w})=\begin{cases}e^{-\mathcal{J}(\mathbf{u},\mathbf{w})}/Z,&Y=1\\ \\ 1-e^{-\mathcal{J}(\mathbf{u},\mathbf{w})}/Z,&Y=0.\end{cases}$ (15)
Consider $E(\gamma)=\mathbb{E}_{(\mathbf{u},\mathbf{w})\sim p_{\gamma}(\mathbf{u},\mathbf{w})}[\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})|Y=1]$, which indicates the expectation of getting an $\epsilon$-optimal solution by using the prior reweighting technique with $\gamma$ under the guidance of $\mathcal{J}$. Define
$F(\gamma)=\frac{\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})\ln(p(\mathbf{w}))]}{\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})]}-\frac{\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)^{c}}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})\ln(p(\mathbf{w}))]}{\mathbb{E}_{\mathbf{u},\mathbf{w}}[\mathbb{I}_{Q(\varepsilon)^{c}}(\mathbf{u},\mathbf{w})p(Y=1|\mathbf{u},\mathbf{w})]},$
then we have the following theorem (please refer to Appendix B for its proof):
Theorem 3.1.
Assume $E(\gamma)$ is a smooth function, then the following hold:
  • If $F(1)<0$, there exists a $\gamma_{-}<1$, s.t., $E(\gamma_{-})>E(1)$;
  • If $F(1)>0$, there exists a $\gamma_{+}>1$, s.t., $E(\gamma_{+})>E(1)$.
Remark: Here $F(\gamma)$ can be interpreted as some kind of difference between "entropies" in $Q(\varepsilon)^{c}$ and $Q(\varepsilon)$. When $F(1)<0$, it means that $Q(\varepsilon)^{c}$ has higher "entropies", implying that the training trajectories are far from optimal. As a result, we may need to flatten the distribution of training trajectories, which corresponds to using the prior reweighting technique with $\gamma<1$. Since this is the most common case in real scenarios, we usually set $\gamma<1$.
@@ 4 Experiments
In this section, we aim to answer the following questions: (1) Can DiffPhyCon present superiority over traditional, supervised learning, and reinforcement learning methods for physical systems control? (2) Does the proposed prior reweighting technique help achieve better control objectives? (3) Could answers to (1) and (2) be generalized to more challenging partial observation or partial control scenarios? To answer these questions, we conduct experiments on three vital and challenging problems: 1D Burgers’ equation, 2D jellyfish movement control, and 2D smoke control problems.
The following state-of-the-art control methods are selected as baselines. For the 1D Burgers’ equation, we use (1) the classical and widely used control algorithm Proportional-Integral-Derivative (PID) [33] interacting with our trained surrogate model of the solver; (2) Supervised Learning method (SL) [24]; RL methods including (3) Soft Actor-Critic (SAC) [16] with offline and surrogate-solver versions; (4) Behaviour Cloning (BC) [49]; and (5) Behavior Proximal Policy Optimization (BPPO) [73]. Specifically, the surrogate-solver version of SAC interacts with our trained surrogate model of the solver, while the offline version only uses given data. BC and BPPO are also in offline versions. For 2D jellyfish movement control, baselines include SL, SAC (offline), SAC (surrogate-solver), BC, BPPO, and an additional classical multi-input multi-output algorithm Model Predictive Control (MPC) [57]. PID is inapplicable to this data-driven task [2]. Detailed descriptions of baselines are provided in Appendix H and Appendix I.
@@
$\mathcal{J}_{\text{actual}}\coloneqq\int_{\Omega}|u(T,x)-u_{d}(x)|^{2}\mathrm{d}x,\ \ \mathcal{J}_{\text{energy}}\coloneqq\int_{[0,T]\times\Omega}|w(t,x)|^{2}\mathrm{d}t\mathrm{d}x.$ (17)
To make the evaluation challenging, we select three different experiment settings that correspond to different real-life scenarios: partial observation, full control (PO-FC), full observation, partial control (FO-PC), and partial observation, partial control (PO-PC), which are elaborated on in Appendix D.2 and illustrated in Appendix C.1. These settings are challenging for classical control methods such as PID since they require capturing the long-range dependencies in the system dynamics. Note that the reported metrics in different settings are not directly comparable. In this experiment, DiffPhyCon uses the guidance conditioning (Section 3.1) to optimize $\mathcal{J}_{\text{actual}}$ and explicit guidance to optimize the energy cost, which is elaborated in Appendix D.3.
@@
To compare the ability of different methods to optimize $\mathcal{J}_{\text{actual}}$ with constrained energy cost $\mathcal{J}_{\text{energy}}$, we compare the Pareto frontiers of different methods in Figure 3. We vary the hyperparameter $\lambda$ to control the tradeoff between $\mathcal{J}_{\text{actual}}$ and the energy cost since most baselines have this hyperparameter. As can be observed in Figure 3, the Pareto frontiers of DiffPhyCon are consistently among the best, achieving the lowest $\mathcal{J}_{\text{actual}}$ for most settings of the energy budget. Although SL performs well in full observation setting (b) where the system dynamics can be more easily predicted, it encounters difficulty in partial observation scenarios (a)(c). The results demonstrate DiffPhyCon’s ability to generate near-optimal control sequences compared to baselines. More visualization results are provided in Appendix C.1. More results of evaluation are presented in Appendix K.1. For efficiency evaluation of training and test phases, please refer to Table 23 in Appendix K.4.
@@
Experiment settings. This task is to control the movement of a flapping jellyfish with two wings in a 2D fluid field where fluid flows at a constant speed. The jellyfish is propelled by the fluid when its wings flap. Its moving speed and efficiency are determined by the mode of flapping. This task is an important source of inspiration for the design of underwater and aerial devices, and its challenges come from complex vortice behavior and fluid-solid coupling dynamics [54, 28]. For this task, fluid dynamics follows the 2D incompressible Navier-Stokes Equation:
@@
subject to Eq. (18) and the boundary condition that the velocity of fluid vanishes near the boundary. The hyperparameter $\zeta$ is set to be 1000. We evaluate in two settings: full observation, where the full state $\mathbf{u}=(\mathbf{v},p)$ is observed; and partial observation, where only pressure is observed. This task is very challenging due to the complicated dynamics of fluid-solid interactions and vortices behaviour [54, 28], especially in the scenario of partial observation where the missing of $\mathbf{v}$ in Eq. (18) restricts the information available to generate well-informed control signals. Details of the experiment are provided in Appendix F.
Open-source Dataset Description. We use the Lily-Pad simulator [66] to generate a dataset describing the movement of jellyfish under the control of its flapping behavior. Statistics about the dataset are listed in Table 2. The feature of this dataset is that it contains a rich variation of vortices behavior and fluid-solid interaction dynamics, determined by violent changes in opening angles of wings and open-close phase ratio. It would serve as an important benchmark for studying complex physical system control problems. Details about the dataset are provided in Appendix E.
@@
Figure 5: Visualization of jellyfish movement and fluid field controlled by DiffPhyCon as in the middle subfigure of Figure 4.
Figure 6: Visualization of smoke density and fluid field dynamics controlled by DiffPhyCon.
Results. Evaluation results are presented in Table 3. It can be seen that our method outperforms the baselines by a large margin in optimizing the control objective $\mathcal{J}$. At a cost of slightly increasing the control cost $R(\mathbf{w})$, control sequences generated by our method achieve a much faster average speed than baselines. In the full observation setting, the control objective achieved by DiffPhyCon is -53.18 lower than the best baseline BPPO, while the average speed exhibits an increment of 172.2 over it. This demonstrates that diffusion models are effective for this challenging control task by performing global optimization of trajectory and control sequences in a generative approach. Comparison between DiffPhyCon-lite and DiffPhyCon reveals that by flattening the prior distribution of control sequences, another significant improvement is further achieved. Even in the more challenging partial observation setting, DiffPhyCon still exhibits substantial advantages over existing methods. This reflects our method has a strong control capability under inadequate information. Configuration of the hyperparameter $\gamma$ in DiffPhyCon and performance with respect to varying $\gamma$ is presented in Figure 17 in Appendix L.1. Details about the hyperparameter $\lambda$ can be found in Table 33 in Appendix L.2.
@@
Figure 4 visualizes generated opening angle curves of different methods on three test jellyfish. Opening angle curves of DiffPhyCon-lite show an obvious fast-close-slow-open shape, which is proven to produce high speed in jellyfish movement [28]. The reason is that the fast closing of wings leads to a thrust from fluid in the early state, resulting in a long-term high speed, and followed by a slow opening to reduce resistance. While this mode of movement appears rarely in the training dataset, DiffPhyCon-lite could generate such control sequences for most test samples. This reflects that diffusion models under guidance are effective in optimizing the control objective. Furthermore, DiffPhyCon makes this mode of movement more aggressive, with sharper change of the opening angle in the beginning and end stage within a period. This provides strong evidence that reweighting the prior distribution of control sequences allows for flexible sampling over a flattened distribution, thus control sequences with low prior but good objectives are more likely to be sampled. The movement and the resulting fluid field of the jellyfish corresponding to the middle subfigure of Figure 4 controlled by DiffPhyCon is illustrated in Figure 5 and more examples are provided in Figure 10 in Appendix C.2. Conversely, opening angles obtained by baselines are inferior. The reason may be that they predict opening angles sequentially and hard to globally optimize the objective with three conflicting terms: average speed, $R(\mathbf{w})$, and $d(\mathbf{w}_{T},\mathbf{w}_{0})$. We further study such myopic failure mode of SAC in Appendix K.3. More comparisons about the variation of the weight $\zeta$ are listed in Table 20 and Table 21 in Appendix K.2. For efficiency evaluation of training and test phases, please refer to Table 24 in Appendix K.4. We also extend this experiment to a high-dimensional control signal setting, where the wings of the jellyfish are assumed to be soft. We find that our method is still competitive with baselines. Details about this evaluation are provided in Appendix M.
4.3 2D Smoke Indirect Control
Experiment settings. This task is to control smoke control in an incompressible fluid environment, following a similar (but more challenging) setup of [22]. Control forces were applied within a $64\times 64$ grid flow field, excluding a semi-enclosed region, to minimize the smoke failing to pass through the top middle exit (seven exits in total). For illustration of our settings, please refer to Figure 14 in Appendix G. This high-dimensional indirect control problem involves managing 2D forces at approximately 1,700 grid points every time step, resulting in about 100,000 control variables across 32 time steps, making it highly challenging.
Table 4: 2D smoke movement control results. Bold font denotes the best model, and underline denotes the second best model.
Method $\mathcal{J}$downarrow
BC 0.3085
BPPO 0.3066
SAC (surrogate-solver) 0.3212
SAC(offline) 0.6503
DiffPhyCon-lite (ours) 0.2324
DiffPhyCon (ours) 0.2254
Results. Evaluation results are presented in Table 4. Our method still has significant advantages over baselines in minimizing the control objective. Furthermore, the prior reweighting technique achieves extra improvement over DiffPhyCon-lite. One test sample of smoke density and fluid field dynamics is illustrated in Figure 6. More visualization results of our method are presented in Figure 11 in Appendix G. These results demonstrate that DiffPhyCon is capable of controlling high dimensional physical systems even when control signals are indirectly applied to the system.